OpenStudy (anonymous):

Given f(x) = 6x^2 - x - 12 and g(x) = 2x - 3, find the function (f/g)(x). Enter your answer as the following example: -10x+1.

OpenStudy (sidsiddhartha):

\[(f/g)(x)=\frac{ f(x) }{ g(x)}\]

OpenStudy (sidsiddhartha):

now just simply divide

OpenStudy (anonymous):

I know that but how do I simplify it??

OpenStudy (sidsiddhartha):

\[=6x^2 - x - 12 =6x^2-(9-8)x-12\] \[=6x^2-9x+8x-12=3x(2x-3)+4(2x-3)\] \[=(3x+4)(2x-3)\] so u have \[f(x)=(2x-3)(3x+4)\] and\[g(x)=2x-3\] \[\frac{ f(x) }{ g(x) }=\frac{ (2x-3)(3x+4) }{ (2x-3)}=(3x+4)\]

OpenStudy (sidsiddhartha):

got this? @MLeighW

OpenStudy (anonymous):

Yes. Thank you!!

OpenStudy (sidsiddhartha):

np :)